OpenStudy (anonymous):

Perform the indicated operations and simplify the result. Leave the answer in factored form.

OpenStudy (anonymous):

\LARGE \frac{\frac{4}{x} + \frac{7}{x^2}}{\frac{16}{x^2} - \frac{49}{x}}

OpenStudy (anonymous):